OpenStudy (anonymous):

Write and then solve for y = f(x) the differential equation for the statement: "The rate of change of y with respect to x is inversely proportional to y^4."

OpenStudy (anonymous):

I beleive it to be dy/dx =1/y^4 ?

OpenStudy (solomonzelman):

Yes, your equation is right.

OpenStudy (solomonzelman):

but, I would say, it is \(\color{#000000}{ \displaystyle \frac{dy}{dx} =\frac{k}{y^4} }\)

OpenStudy (anonymous):

to simulate some constant

OpenStudy (solomonzelman):

yes, because if you said that y is proportional to x, then y=kx. In other words, there is some constant k such that x•k=y.

OpenStudy (solomonzelman):

The simplest here is to separate the variables.

OpenStudy (anonymous):

so then I would get that y^5/5 =kx

OpenStudy (anonymous):

+c

OpenStudy (anonymous):

integrate 1/y^4. so \[y^4dy=kdx , \int\limits_{}^{}y^4dy=\int\limits_{}^{}kdx\] \[\frac{ y^5 }{ 5 }=kx+c\]

OpenStudy (anonymous):

I think thats adequite, isolating y would be nice but i dont think its necessary, the equation is simple in this form

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \frac{dy}{dx} =\frac{k}{y^4} \quad \Longrightarrow \quad y^4\frac{dy}{dx} =k }\) I like claiming that you don't just "integrate", rather you integrate with respect to a variable. (and then cancel differentials) \(\color{#000000}{ \displaystyle \color{red}{\int} y^4\frac{dy}{dx} \color{red}{dx} =\color{red}{\int} k \color{red}{dx} }\) \(\color{#000000}{ \displaystyle \int y^4\frac{dy}{\cancel{dx}} \cancel{dx}=\int k dx }\) \(\color{#000000}{ \displaystyle \int y^4~dy=\int k~dx }\) and then integrate both sides, BUT DON'T FORGET ABOUT INTEGRATION CONSTANT C. Or, if you prefer, you can just go ahead and "integrate" without canceling differentials as I showed. That is, \(\color{#000000}{ \displaystyle y^4~= k~ \quad \Longrightarrow \quad \int y^4~dy=\int k~dx}\) and either way (when done correctly) you end up with the same result.

OpenStudy (solomonzelman):

oh excuse me let me fix my last equation

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle y^4\frac{dy}{dx}=k\quad\Longrightarrow\quad y^4{\tiny~}dy=k{\tiny~}dx \quad\Longrightarrow\quad \int y^4{\tiny~}dy=\int k{\tiny~}dx }\)

OpenStudy (anonymous):

i usuallyt look at it as having to seperate dy/dx by multiplying dx out ect

OpenStudy (solomonzelman):

Like I wrote just now?

OpenStudy (anonymous):

Well i would conuse myself with thinking for example, that I have dy=45y^2/2x dx and when i went to divide the 45y^2 i would think i could not becuase i would effectively be making dy/45y^2. usually i made this error when tired but once i ingrained it into myself to disassociate the dx or dy with any other things i do in the equation it worked out

OpenStudy (solomonzelman):

Ok)

OpenStudy (anonymous):

But yeah It seemed dumb once i made sure but it was a common mistake for me for a while

OpenStudy (solomonzelman):

So, basically, when separating variables, all you want to do is to rearrange your original equation (without preforming incorrect operations, and when possible of course), isuch that it is something like \(\color{#000000}{ \displaystyle f(y)~dy~=~f(x)~dx }\)

OpenStudy (solomonzelman):

and from there you can simply integrate both sides.

OpenStudy (anonymous):

Yeah i understand all of that I was just sharing a personal struggle i had

OpenStudy (solomonzelman):

Alright.

OpenStudy (solomonzelman):

so, in the original problem we left off, \(\color{#000000}{ \displaystyle \int y^4~dy=\int k~dx }\)

OpenStudy (solomonzelman):

now, all you need to do is to integrate, but don't forget arbitrary constants.

OpenStudy (anonymous):

oh yeah i alreasy finished the problem if you look back

OpenStudy (solomonzelman):

yes, you say, y^5/5=kx

OpenStudy (solomonzelman):

and that is incomplete

OpenStudy (anonymous):

I made the +c adendam

OpenStudy (solomonzelman):

Oh, ok.

OpenStudy (solomonzelman):

Missed that; apologize.

OpenStudy (anonymous):

NP I appreciate the help!

OpenStudy (solomonzelman):

So, your final is, \(\color{#000000}{ \displaystyle \frac{1}{5}y^5=kx+C}\)

OpenStudy (solomonzelman):

you are welcome, if I helped:)

OpenStudy (anonymous):

you did!

OpenStudy (solomonzelman):

Just curious is this calculus I, II or a differential equatiosn class?

OpenStudy (anonymous):

calc 1 one of the easier ones on the test just had not worked with a word problem in a while so i was unsure of myself

OpenStudy (solomonzelman):

Yeah, I am not a great fan of them either. Especially if it is poorly worded, which it would usually be the case as it turns out....

OpenStudy (anonymous):

yeah these practice tests can be brutal though sometimes I had a pretty bad run in with the area of a revolved object of which half of the instructions indicated a square and the other half a circle spent like 90 min on openstudy nobody could figure it out XD

OpenStudy (solomonzelman):

still have problem with those, or you solved them?

OpenStudy (anonymous):

Oh no I am great at that those come easy to me as im cerified in 2 cad softwares so it was like second nature, but the wording was completely incorrect

OpenStudy (solomonzelman):

yes, fortunatelly, it is readable mostly.

OpenStudy (solomonzelman):

in any case, won't disturb.... good luck with calc!